OpenStudy (anonymous):

An airplane leaves an airport and flies due west 150 miles and then 170 miles in the direction S 49.17°W. How far is the plane from the airport (round to the nearest mile)?

OpenStudy (anonymous):

I got ~300 mi?

OpenStudy (anonymous):

Its not right is it?

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

Anyone?...

OpenStudy (phi):

you need to use trigonometry to find x and y then pythagoras to find d. what did you get for x? for Y?

OpenStudy (anonymous):

i used the law of cosines formula?

OpenStudy (anonymous):

b^2=a^2+c^2-2ac(cosB)

OpenStudy (anonymous):

and then found the sqrt to ge tb which was about 300...my answer

OpenStudy (phi):

yes, that will work if you use angle = 90+49.17